# Learning New Tasks from a Few Examples with Soft-Label Prototypes

### v3

@@ 5 Tasks and datasets
We use the same tasks and datasets as in Bansal et al. (2019). These cover a variety of text classification tasks with a large number of classes, and include: (a) Entity typing – the CoNLL-2003 Sang and De Meulder (2003) and MIT-Restaurant Liu et al. (2013) datasets; (b) Review rating classification – review ratings from Amazon Reviews Blitzer et al. (2007) with a three-way classification; (c) Text classification – scraped social media data from crowdflower [^2] https://www.figure-eight.com/data-for-everyone/ comprising sentiment and emotion classification in a range of domains, as well as political bias detection; and (d) Natural language inference in the scientific domain – the SciTail dataset Khot et al. (2018). We randomly sample $1000$ instances from the emotion task in (c) and use them as our held-out validation set for hyperparameter tuning.
We use the same data splits as Bansal et al. (2019), which are publicly available. [^3] https://github.com/iesl/leopard/ At evaluation time, Bansal et al. (2019) sample an N-way k-shot training (support set) subset $\forall\;k\in\{4,8,16\}$ for few-shot fine-tuning and then evaluate model performance on the test set of each respective task. As performance is dependent on the data used for fine-tuning, they sample $10$ different subsets per task, and report the averaged performance across subsets on the test set. We utilise the same subsets as Bansal et al. (2019) for fine-tuning our models too. Unlike Bansal et al. (2019), however, we do not utilise auxiliary training data (GLUE; Wang et al. (2018)), and only train our models on the few-shot training subsets provided. This makes our setting more challenging but one which is testing a more realistic few-shot setting that does not rely on additional data and labels. While we are not directly comparable to Bansal et al. (2019), to facilitate such a comparison, we present an analysis where we also utilise GLUE data (Section 9.2).
Similar to Bansal et al. (2019), we use mean accuracy and standard deviation to evaluate performance. For entity typing tasks, we obtain token-level encodings for classification.
@@ 6 Baselines
Our aim is to determine how well our soft-label prototype models perform compared to other classifiers given the same few-shot training data. Again, our focus is on performing an extreme few-shot learning scenario where no auxiliary data is available, and so we choose such baseline comparisons accordingly.
1-Nearest Neighbour Classifier
Our first baseline is a 1-Nearest Neighbour classifier for each task. This is motivated by the fact that test time predictions using soft-label prototypes depend on the soft-label functions of the closest class centroids to a test point $x$ (Section 3.2). We first encode all data points using BERT, and then assign test instances the class of their closest training example in this embedding space based on Euclidean distance.
Fine-tuned BERT
We use the BERT baseline reported in Bansal et al. (2019), which is fine-tuned (all layers) on the training set of each task and is trained using the same experimental/few-shot setting as ours. We refer to this model as BERTfine-tuned.
@@
We use Prototypical networks (ProtoNets) Snell et al. (2017) as another baseline to evaluate generalisation within a few-shot meta-learning setup. ProtoNet is a metric-based meta-learning method that uses an embedding function $f_{\theta}$ to encode training samples and compute a high-dimensional vector $\mu_{c}$ that is the arithmetic mean of the training data points of class $c$. It then uses a distance function $d$ to compute the similarity between a test instance $x$ and the mean vector of each class, and get the class probability distribution as:

### v4

@@ 4 Experimental settings and datasets
We use the same test datasets and evaluation setting as Bansal et al. (2020a) for both the high-resource and low-resource settings. These cover a variety of text classification tasks: (a) Entity typing – the CoNLL-2003 Tjong Kim Sang and De Meulder (2003) and MIT-Restaurant Liu et al. (2013) datasets; (b) Review rating classification – review ratings from Amazon Reviews Blitzer et al. (2007) with a three-way classification; (c) Text classification – scraped social media data from crowdflower comprising sentiment and emotion classification in a range of domains, as well as political bias detection; and (d) Natural language inference in the scientific domain – the SciTail dataset Khot et al. (2018). We use the same data splits, which are publicly available. During evaluation, Bansal et al. (2020a) fine-tune their models using a small few-shot (support) training set per test task (using a $k$-shot setting of $4,8,16$ examples per class), and then evaluate performance on each task’s dedicated, unseen test set. As model performance can be affected by the $k$ examples chosen for training/fine-tuning, for each task and for every k, they sample $10$ few-shot training sets and report the mean and standard deviation, which we also adopt in our experiments.
@@ 5 Baselines
Our aim is to determine how well our models – DeepSLP and MetaSLP – perform in the low and high-resource setting respectively, compared to strong baselines when given the same few-shot adaptation sets. Our focus is on (a) evaluating in an extreme few-shot learning scenario where no auxiliary data is available (using DeepSLP), and (b) evaluating in a high-resource setting when additional (auxiliary) training data is available (MetaSLP). We use BERT Devlin et al. (2019) as our text encoder throughout to facilitate model comparisons. We report DeepSLP and MetaSLPREPTILE (i.e., using Reptile as our meta-learning algorithm) in our main table of results (given their effectiveness), and present additional baselines as well as hyperparameters and training details in Appendix A.6.
We use BERT Devlin et al. (2019) as our encoder as it is a text encoder that allows us to get passage-level encodings, it is computationally light-weight compared to decoder-based LLMs such as Llama Touvron et al. (2023) and GPT OpenAI (2024) (and we can carry out full fine-tuning) and, crucially, it does not suffer from data contamination as, in contrast to the more recent LLMs, it has not been pre-trained on our test data [^3] https://github.com/iesl/leopard/tree/master/data/json.
5.1 Low-resource setting baselines
BERTfine-tuned
We use BERTfine-tuned reported in Bansal et al. (2020a), which is fine-tuned (all layers) on the few-shot training set of each test task.
LORABERT
LORA Hu et al. (2022) decomposes a fine-tuned weight matrix to two low-rank matrices, which – when multiplied and added to the original weights – reproduce the fine-tuned weights. This is advantageous as, instead of fine-tuning all parameters, we fine-tune these two matrices with a low computational cost, as they are much smaller individually compared to fully fine-tuned weights.
constraintSLP
We use constraintSLP as a baseline to evaluate the effectiveness of soft-label prototypes that are based on linear constraints.
5.2 High-resource setting baselines
Reptile
We train a meta-learning Reptile Nichol et al. (2018) model on our auxiliary data and use it as another baseline. This allows us to directly assess the added advantage of utilising SLPs in MetaSLPREPTILE.
@@ 5.2 High-resource setting baselines
We use ProtoNet Snell et al. (2017) as another baseline for both the high and low-resource settings. ProtoNets use Euclidean distance as a measure of similarity between points and clusters, which is similar to DeepSLP and MetaSLP that assign test points to the closest line.
LEOPARD
Bansal et al. (2020a) present LEOPARD, a meta-learning algorithm that achieves the best performance across most test tasks for entity typing, ratings classification and text classification, and which we also use.
We do not include HSMLMT Bansal et al. (2020b) as a baseline as it is pretrained on semi-supervised meta-training tasks in addition to supervised learning and therefore it is not directly comparable.
Category (Classes) Shot LORABERT BERTfine-tuned* DeepSLP LEOPARD* Reptile MetaSLPREPTILE
Political Bias (2) 4 52.75 ± 4.33 54.57 ± 5.02 53.251 ± 4.042 60.49 ± 6.66 58.82 ± 4.31 60.96 ± 6.13
8 53.66 ± 4.25 56.15 ± 3.75 58.209 ± 5.198 61.74 ± 6.73 59.43 ± 3.79 63.65 ± 4.57
16 59.21 ± 2.27 60.96 ± 4.25 61.479 ± 2.974 65.08 ± 2.14 62.21 ± 0.72 66.05 ± 1.57
Emotion (13) 4 7.56 ± 2.93 09.20 ± 3.22 9.076 ± 1.108 11.71 ± 2.16 11.65 ± 3.21 11.94 ± 1.95
8 9.02 ± 2.36 08.21 ± 2.12 8.041 ± 2.797 12.90 ± 1.63 10.56 ± 2.85 13.42 ± 1.46
16 10.29 ± 1.67 13.43 ± 2.51 10.919 ± 1.615 13.38 ± 2.20 11.62 ± 3.11 14.03 ± 2.35
Sentiment Books (2) 4 51.27 ± 2.75 54.81 ± 3.75 58.67 ± 4.753 82.54 ± 1.33 76.95 ± 1.03 83.22 ± 0.95
8 58.16 ± 3.3 53.54 ± 5.17 64.78 ± 2.615 83.03 ± 1.28 77.49 ± 1.08 83.8 ± 0.8
16 59.16 ± 2.59 65.56 ± 4.12 67.453 ± 3.085 83.33 ± 0.79 77.88 ± 0.56 83.8 ± 1.59
Rating DVD (3) 4 31.65 ± 4.91 32.22 ± 08.72 39.566 ± 5.086 49.76 ± 9.80 45.91 ± 9.85 45.2 ± 8.91
8 37.69 ± 3.16 36.35 ± 12.50 38.788 ± 4.449 53.28 ± 4.66 47.23 ± 9.22 58.38 ± 2.9
16 38.63 ± 5.52 42.79 ± 10.18 40.53 ± 4.375 53.52 ± 4.77 48.49 ± 8.88 57.41 ± 4.71
Rating Electronics (3) 4 31.66 ± 2.94 39.27 ± 10.15 39.977 ± 5.959 51.71 ± 7.20 44.47 ± 8.25 45.34 ± 7.22
8 38.72 ± 5.95 28.74 ± 08.22 41.926 ± 3.985 54.78 ± 6.48 49.1 ± 6.81 55.10 ± 5.12
16 39.15 ± 6.6 45.48 ± 06.13 44.917 ± 3.164 58.69 ± 2.41 50.68 ± 6.8 59.47 ± 2.29
Rating Kitchen (3) 4 36.63 ± 4.68 34.76 ± 11.20 39.624 ± 6.787 50.21 ± 09.63 45.38 ± 10.96 45.20 ± 8.78
8 39.69 ± 6.22 34.49 ± 08.72 41.081 ± 6.777 53.72 ± 10.31 46.71 ± 9.84 54.53 ± 9.9
16 38.17 ± 7.14 47.94 ± 08.28 45.801 ± 4.562 57.00 ± 08.69 52.87 ± 9.52 58.94 ± 7.58
Political Audience (2) 4 49.75 ± 1.03 51.02 ± 1.72 51.741 ± 2.827 52.60 ± 3.51 52.45 ± 4.26 54.1 ± 3.66
8 54.05 ± 2.54 52.80 ± 2.72 54.506 ± 3.274 54.31 ± 3.95 52.87 ± 4.31 56.01 ± 3.65
16 55.39 ± 3.66 58.45 ± 4.98 56.956 ± 3.045 57.71 ± 3.52 55.6 ± 1.85 58.57 ± 2.04
Sentiment Kitchen (2) 4 53.02 ± 1.54 56.93 ± 7.10 60.76 ± 4.426 78.35 ± 18.36 69.81 ± 14.58 81.96 ± 3.73
8 55.54 ± 3.47 57.13 ± 6.60 65.733 ± 3.198 84.88 ± 1.12 75.76 ± 1.13 83.33 ± 1.99
16 58.59 ± 4.83 68.88 ± 3.39 69.18 ± 2.589 85.27 ± 1.31 76.41 ± 0.66 84.33 ± 1.81
Disaster (2) 4 56.02 ± 6.35 55.73 ± 10.29 54.252 ± 9.843 51.45 ± 4.25 49.76 ± 4.73 55.03 ± 8.73
8 57.46 ± 6.9 56.31 ± 09.57 61.3 ± 7.961 55.96 ± 3.58 52.17 ± 5.17 57.77 ± 6.40
16 65.79 ± 2.03 64.52 ± 08.93 69.28 ± 2.358 61.32 ± 2.83 55.37 ± 4.53 65.18 ± 4.41
Airline (3) 4 24.36 ± 5.42 42.76 ± 13.50 50.987 ± 4.936 54.95 ± 11.81 57.11 ± 14.16 57.39 ± 7.83
8 52.31 ± 7.89 38.00 ± 17.06 55.209 ± 6.049 61.44 ± 03.90 64.37 ± 3.49 65.67 ± 4.82
16 54.1 ± 8.57 58.01 ± 08.23 60.247 ± 4.577 62.15 ± 05.56 66.31 ± 2.55 69.48 ± 2.06
Rating Books (3) 4 34.69 ± 2.12 39.42 ± 07.22 42.116 ± 4.725 54.92 ± 6.18 56.57 ± 8.17 55.79 ± 5.61
8 39.36 ± 6.33 39.55 ± 10.01 42.156 ± 4.608 59.16 ± 4.13 57.33 ± 7.63 65.74 ± 5.58
16 41.23 ± 5.32 43.08 ± 11.78 46.513 ± 3.036 61.02 ± 4.19 63.26 ± 3.59 67.87 ± 3.45
Political Message (9) 4 12.16 ± 1.46 15.64 ± 2.73 14.421 ± 1.095 15.69 ± 1.57 14.58 ± 1.78 18.84 ± 1.82
8 15.71 ± 2.04 13.38 ± 1.74 16.919 ± 1.756 18.02 ± 2.32 15.13 ± 2.16 20.09 ± 2.71
16 15.53 ± 2.55 20.67 ± 3.89 18.319 ± 1.74 18.07 ± 2.41 16.38 ± 2.15 23.22 ± 1.17
Sentiment DVD (2) 4 50.77 ± 0.78 54.98 ± 3.96 55.003 ± 2.936 80.32 ± 1.02 72.03 ± 11.61 80.97 ± 1.21
8 52.24 ± 1.54 55.63 ± 4.34 57.527 ± 3.562 80.85 ± 1.23 75.79 ± 1.62 81.85 ± 1.79
16 52.6 ± 2.09 58.69 ± 6.08 60.76 ± 2.944 81.25 ± 1.41 76.69 ± 0.8 83.48 ± 1.01
Scitail (2) 4 43.36 ± 4.74 58.53 ± 09.74 54.101 ± 3.759 69.50 ± 9.56 59.13 ± 10.58 53.48 ± 5.59
8 54.29 ± 5.25 57.93 ± 10.70 56.341 ± 5.786 75.00 ± 2.42 62.63 ± 10.85 60.79 ± 4.6
16 52.68 ± 3.0 65.66 ± 06.82 59.692 ± 4.227 77.03 ± 1.82 68.03 ± 1.57 61.67 ± 3.61
Restaurant (8) 4 10.56 ± 1.36 49.37 ± 4.28 47.634 ± 5.237 49.84 ± 3.31 13.37 ± 2.25 27.00 ± 2.61
8 20.92 ± 2.4 49.38 ± 7.76 55.912 ± 4.494 62.99 ± 3.28 16.83 ± 3.42 35.66 ± 2.39
16 29.37 ± 4.05 69.24 ± 3.68 61.716 ± 2.208 70.44 ± 2.89 16.0 ± 3.44 37.20 ± 2.68
CoNLL (4) 4 21.48 ± 2.71 50.44 ± 08.57 52.724 ± 5.84 54.16 ± 6.32 31.31 ± 5.32 40.79 ± 3.40
8 29.84 ± 3.28 50.06 ± 11.30 60.374 ± 3.731 67.38 ± 4.33 33.17 ± 5.1 41.25 ± 5.21
16 37.18 ± 3.32 74.47 ± 03.10 67.496 ± 4.551 76.37 ± 3.08 34.04 ± 3.59 45.96 ± 4.75
Table 1: Classification performance (accuracy) of our methods (DeepSLP and MetaSLP) and baselines. * refers to the baselines as reported in Bansal et al. (2020a). The best performing models for each setting (without and with auxiliary data) are highlighted in gray and green respectively. Double lines group similar tasks together: the first set contains intent classification tasks, the second focuses on natural language inference, and the last contains entity typing tasks.
6 Results and Discussion
Due to space restrictions, we present and discuss our best models in Table 1. All other baselines and results are discussed in Appendix A.7.
Low-resource setting
In the low-resource setting with no auxiliary data (left side of Table 1), DeepSLP outperforms all baselines, including BERTfine-tuned in $31/48$ tasks and LORABERT in $45/48$ tasks, achieving a new state-of-the-art result. Our results demonstrate the usefulness of soft-label prototypes and their superiority over strong baselines (i.e., LLM fine-tuning and LORA / low-rank adaptation) in the low-resource setting.
Unlike BERTfine-tuned, DeepSLP and LORABERT do not fine-tune the encoder. Specifically, we only need to fine-tune $1500-10K$ parameters (based on the number of classes) for each line with two soft-label prototypes for DeepSLP, compared to $50K-100K$ parameters for LORABERT with $rank=2$, and $>10^{8}$ parameters for BERTfine-tuned. We also note that DeepSLP is lightweight and does not require a GPU. LORABERT mostly achieves accuracies within $90\%$ of BERTfine-tuned, in line with previous work Hu et al. (2022); Dettmers et al. (2023) (even outperforming BERTfine-tuned in $15/48$ tasks), except for entity-typing tasks where LORABERT struggles to generalise and achieves substantially lower performance compared to BERTfine-tuned.
On the other hand, constraintSLP, a simpler variant of DeepSLP (see Appendix A.7, Table 5 for results) is one of the lower performing baselines, together with ProtoNet. We find that constraintSLP exhibits a substantial weakness (see further details in Theorem A.2, Appendix A.3): given Euclidean distance, constraintSLP does not always select the nearest class centroid to a test point. This violates our inductive bias that points located closest to a class centroid are assigned to that class. If we consider the case where $N=2$, constraintSLP essentially acts as a 1-NN with soft labels trivially at $[1,0]$ and $[0,1]$, with class centroids acting as the nearest neighbour. However, when generalising beyond this setting, the model’s stability is affected. constraintSLP optimises soft labels using the geometric properties of a line and does not consider each (training) data point individually – the soft labels produced by SLP are constants. DeepSLP, on the other hand, learns from training data and produces soft labels as a function of the input; therefore, it has the ability to output soft labels based on the location of an input (test) point (with the location of the prototypes being fixed).
High-resource setting
MetaSLPREPTILE has the highest performance overall in text classification and entailment tasks (Tasks 1-14), with the best accuracy in $33/42$ tasks/settings. LEOPARD, on the other hand, achieves the highest score in only $8/42$. Interestingly, we find that all models in the high-resource setting have lower performance for Disaster compared to the models in the low-resource setting. We surmise this to be due to the auxiliary data and the fact that the meta-training distribution differs substantially from the test distribution.
For entity typing tasks (CoNLL and Restaurant), LEOPARD outperforms all models, with MetaSLPREPTILE and Reptile performing comparatively poorly, even outperformed by the low-resource methods (DeepSLP and BERTfine-tuned). It should be noted that there seems to be little benefit of meta-learning with auxiliary data when tackling entity typing tasks, even for LEOPARD, as the difference between LEOPARD and BERTfine-tuned is not substantial. We surmise this to be due to the fact that the meta-training distribution (i.e., GLUE tasks) is different from the test distribution for entity typing tasks which degrades performance for the test tasks. Note that we do not meta-train the entire model (only top 4 layers for Reptile and MetaSLPREPTILE), unlike LEOPARD.
Overall, MetaSLPREPTILE outperforms all models and baselines, including Reptile in $42/48$ tasks and LEOPARD in $34/48$ tasks. Specifically, MetaSLPREPTILE consistently outperforms Reptile, demonstrating the effectiveness of our approach over its meta-learning variant (i.e., Reptile) that does not use SLPs. In Appendix A.8 we present detailed analyses of DeepSLP and show that it displays several desirable properties of ensemble methods which drive its performance, in addition to it being a computationally efficient approach that only utilises a small number of parameters.
7 Conclusion and future work
We presented a novel few-shot learning paradigm that is based on soft-label prototypes capturing the simultaneous membership of data points over several classes, and demonstrated its effectiveness in low and high-resource settings. We evaluated our approach on $48$ different tasks / settings and showed that it outperforms a range of strong baselines. In the future, we plan to use meta-learning algorithms such as PACMAML Ding et al. (2021) and Bayesian MAML Kim et al. (2018) that relax assumptions with respect to train–test set distributions and thus alleviate this current limitation in our work.
8 Ethics
To the best of our knowledge, there are no ethical concerns involved in this research. We conduct our work using publicly available English datasets and tasks, and models pre-trained on English text. Our results may not generalise to other languages. To facilitate further research in the field, we release our source code and models.
Appendix A Appendix
A.1 Deriving soft-label prototypes using constraintSLP
Finding lines connecting all centroids
In Figure 3(a), we present an example set of lines connecting all class centroids. For further details on recursive regression, we refer the reader to Sucholutsky et al. (2021).
Deriving soft-label prototypes by optimising for linear constraints
Example soft-label prototypes which are “set” at the ends of each line are shown in Figure 3(b).
Classification with constraintSLP: A toy example
Figure 4 presents an example classification with soft-label prototypes. Given the class centroids for blue, green and yellow are located at $(0,0)$, $(1.5,0)$ and $(3,0)$ respectively, two soft-label prototypes are defined by a line connecting yellow and blue, and are thus located at $(3,0)$ and $(0,0)$ respectively. The soft labels in Figure 4(a) contain the per-class probability distribution derived by the constraintSLP method; for example, $p(x=blue)=0.6$ and $p(x=green)=0.4$ for the left prototype, and $p(x=green)=0.4$ and $p(x=yellow)=0.6$ for the right prototype. When a new test instance $x$ located at $(1.5,0.8)$ is presented, we make predictions as follows: we find the nearest line to $x$ and consider its distance from the two prototypes at the ends of the line and multiply the class distribution of each prototype by the inverse distance as per Eq. 1.
Figure 4: Classification example with constraintSLP (figure from Sucholutsky and Schonlau (2021)).
A.2 Meta-learning
For encoder-based models, meta-learning has emerged as a viable methodology for few-shot learning. In the meta-learning paradigm, the training and test sets, referred to as $\mathcal{D}$meta-train and $\mathcal{D}$meta-test, are split into episodes. Each episode encompasses a task $\mathcal{T}$i and consists of a support set $\mathcal{D}^{(i)}$support and a query set $\mathcal{D}^{(i)}$query. Meta-learning algorithms initially fit the model on the support set of the episode (inner-loop optimisation) and then achieve generalisation across episodes by optimising performance on the query sets of the episodes (outer-loop optimisation). For evaluation, the model is first fine-tuned on the support set and then evaluated on the query set for each task $\mathcal{T}$i $\in$ $\mathcal{D}$meta-test. We describe the process algorithmically in Algorithm 2 and describe the MetaUpdate process for different algorithms subsequently.
1 $\alpha,\beta\leftarrow$ learning rates
2 Sample batch of tasks $\{T_{i}\}\sim p(T)$
3 Initialise $\mathbf{\theta^{\prime}_{i}}\leftarrow\theta$
4 for $T_{i}\sim p(T)$ do
5       Partition $T_{i}$ into $D^{s}_{i}$ and $D^{q}_{i}$
6       $\mathbf{\theta^{\prime}_{i}}\leftarrow\mathbf{\theta}-\alpha\nabla_{\theta}\mathcal{L}^{s}_{\mathcal{D}_{i}}(f_{\theta})$ for $k$ steps
7
8 end for
[…]
Algorithm 2 Meta-learning
Model Agnostic Meta-Learning
MAML Finn et al. (2017) is an optimisation-based meta-learning approach which incorporates generalisability across tasks in its cost function. The task loss $\mathcal{L}^{q}_{\mathcal{T}_{i}}$ is computed on the query examples in each episode, using this task-specific model. The initial model parameters $\theta$ are then updated so as to minimize the sum of the losses of all tasks in a batch, leading to improved generalisation across tasks. The MetaUpdate step is thus defined as
$\mathbf{\theta}\leftarrow\mathbf{\theta}-\beta\>\nabla_{\theta}\sum_{\mathcal{T}_{i}\sim p(\mathcal{T})}\mathcal{L}^{q}_{\mathcal{D}_{i}}(f_{\theta_{i}^{\prime}})$
Note that the MetaUpdate expression calculates the gradients of each $\theta_{i}$ with respect to $\theta$, thus necessitating the computation of second-order gradients. To ease computation, we use a first-order approximation of MAML (FOMAML) wherein the gradients of each $\theta_{i}$ are calculated with respect to $\theta_{i}$ and reduce the MetaUpdate term to
$\mathbf{\theta}\leftarrow\mathbf{\theta}-\beta\>\nabla_{\theta_{i}}\sum_{\mathcal{T}_{i}\sim p(\mathcal{T})}\mathcal{L}^{q}_{\mathcal{D}_{i}}(f_{\theta_{i}^{\prime}})$
LEOPARD
Bansal et al. (2020a) employ meta-learning for diverse NLP tasks in an approach inspired from MAML which integrates a text encoder model with a meta-learned parameter generator to tailor task-specific initialisations for the classification head. Their inner-loop update learns the parameter generator for the task, adapts task-specific model parameters and the MetaUpdate step adapts model parameters as done in MAML. They show that their method, LEOPARD, outperforms multi-task trained models as well as a range of other meta-learning methods.
Reptile
This meta-learning algorithm, introduced by Nichol et al. (2018), is computationally simple compared to MAML and LEOPARD - the MetaUpdate step simply moves the model parameters towards inner-loop fine-tuned model parameters, thus assuming the form:
$\theta\leftarrow\theta+\beta\frac{1}{|\{T_{i}\}|}\sum_{T_{i}\sim p(T)}(\theta_{i}-\theta)$
Despite it’s simplicity, it reports strong performance on a variety of few-shot learning tasks Dou et al. (2019).
Prototypical Networks
@@ 5       Partition $T_{i}$ into $D^{s}_{i}$ and $D^{q}_{i}$
Unlike optimisation-driven meta-learning methods, Prototypical Networks Snell et al. (2017) is a metric-based meta-learning method that uses an embedding function $f_{\theta}$ to encode training support samples and compute a high-dimensional vector $\mu_{c}$ that is the arithmetic mean of the training data points of class $c$. It then uses a distance function $d$ to compute the similarity between a query instance $x$ and the mean vector of each class to get the class distribution as: